OpenStudy (anonymous):

the function f(x)=tan x is defined in such a way that f-1 is a function. what can be the domain of f(x)

OpenStudy (anonymous):

restrict is severely

OpenStudy (anonymous):

usual way is to restrict the domain to \((-\frac{\pi}{2},\frac{\pi}{2})\)

OpenStudy (mathmale):

lyubas, do you mean\[f ^{-1}(x)?\]

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

probably not

OpenStudy (anonymous):

how did u gt tht answer? @satellite73

OpenStudy (anonymous):

in order to have an inverse that is a function, you have to restrict the domain of tangent so that it is a one to one function. then it will have an inverse you are not restricting the domain of \(\tan^{-1}(x)\) that domain is all real numbers

OpenStudy (anonymous):

on that interval tangent is strictly increasing and so it a one to one function

OpenStudy (anonymous):

how do u no its a one to one function tho?

OpenStudy (anonymous):

there r no y's in this problem to look at

OpenStudy (anonymous):

and how do u restrict the domain of a function?

OpenStudy (mathmale):

I'd suggest that you, lyubas, draw a couple of periods of the tangent function. This alone would show you why you need to restrict the domain of the tangent function before you try to find the inverse function \[\tan ^{-1}x.\]

OpenStudy (anonymous):

it is your function, you can restrict the domain any way you like

OpenStudy (mathmale):

lyubas: in your shoes I would do an Internet search for "tangent function". this would almost surely turn up graphs of this tangent function and help you towards understanding what restrictions on the domain are necessary. Hint: tan pi/2 is undefined.

OpenStudy (anonymous):

ahh ok thnks

OpenStudy (anonymous):

can u help me wit more questions or u have to go?mathmale

OpenStudy (mathmale):

Most commonly, the graph of the tangent function is depicted as beginning just after -pi/2 and ending just before pi/2.

OpenStudy (mathmale):

so, are you comfortable with this problem, lyubas?

OpenStudy (anonymous):

im good il do more research later to understand it betta thnks for caring :)

OpenStudy (mathmale):

Great! Please post your next question in the usual way.

OpenStudy (anonymous):

kk wil do